OpenStudy (anonymous):

-2(a-8)+6a=36 What is the value of a?

pooja195 (pooja195):

−2(a−8)+6a=36 (−2)(a)+(−2)(−8)+6a=36(Distribute) −2a+16+6a=36 (−2a+6a)+(16)=36(Combine Like Terms) 4a+16=36 4a+16=36 Subtract 16 from both sides 36-16= ?

OpenStudy (anonymous):

20

pooja195 (pooja195):

4a=20 divide both sides by 4 20/4=?

OpenStudy (anonymous):

5

pooja195 (pooja195):

Good so a=5 :=)

OpenStudy (anonymous):

thank you

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)